OpenStudy (anonymous):

Calculate the density of a substance in pen- nyweights per peck that has a volume of 0.333 hogsheads and a mass of 49.62 kg. 1 pennyweight = 1.55 grams 1 hogshead = 238.5 liters 1 peck = 9.091 liters. Answer in units of pw/pk

OpenStudy (anonymous):

Alright we know that \[D=\frac{M}{V}\] So I would do this in two different factor lable solutions. In order to find the M value and the V value. \[49.62 kg (\frac{1000 g}{1 kg})(\frac{1 pyw}{1.55g})=3.201x10^4 pyw\] now solve for the V value. \[0.333 hog(\frac{238.5 L}{1 hog})(\frac{1 peck}{9.091 L})=8.74 peck\] now we will put into the density formula and solve for D \[\frac{3.201x10^4}{8.74 peck}=D\] \[D=3662 \frac{pennywight}{peck}\]

OpenStudy (anonymous):

THANK YOU SO MUCH <3

OpenStudy (anonymous):

No problem I would check sig figs if thats an issue you are dealing with in class